Istituto ed Osservatorio Astronomico Accepted Observation Proposals

1997-1993

(written Jul 2 1997)

[1] G. Peres, R. Betta, A. Ciaravella, S. Orlando, F. Reale & D. Spadaro, “Coronal Loops and EUV Cool Loops Rooted in Sunspots,” 1997, Proposal selected by NASA for SOHO (AO-1) Observation Program.

[2] F. Reale, R. Betta, A. Ciaravella, S. Orlando, G. Peres, D. Spadaro & R. Ventura, “Diagnostics of Heating Episodes in Active Regions,” 1997, Proposal selected by NASA for SOHO (AO-1) Observation Program.

[3] A. Maggio, F. Favata, N. Brickhouse & A.K. Dupree, “Coronal Spectra of Metal-Rich Late-Type Stars,” 1996, Accepted Proposal for ASCA Observation (AO-5) .

[4] S. Sciortino, F. Reale, R. Rosner, S. Serio & M. Barbera, “The coronae of low-mass active stars,” 1996, Proposal selected for SAX Core Program.

[5] F. Favata, A. Maggio, N. Brickhouse, A. Dupree & G. Tagliaferri, “X-ray Spectroscopy of the W Uma system 44 Boo,” 1996, Proposal selected for SAX Core Program.

[6] F. Favata, G. Micela, R. Mewe & R. Pallavicini, “SAX Observations of Capella, VY ARI and lambda And,” 1996, Proposal selected for SAX Core Program.

[7] R. Pallavicini, G. Micela, M. S. White, R. Mewe, M. Kuerster, G. Tagliaferri, A. Ortolani, A. Maggio, S. Randich & L. Pasquini, “Probing the corona of the young star AB Dor,” 1996, Proposal selected for SAX Core Program.

[8] R. Pallavicini, A. Maggio, R. Mewe, A. Ortolani, L. Pasquini, S. Randich, S. Sciortino, G. Tagliaferri & S. M. White, “Coronal Diagnostics of HR 1099, and UX Ari,” 1996, Proposal selected for SAX Core Program.

[9] A. Maggio, G. Peres, S. Sciortino & R. Rosner, “Thermal Structure of Late-type Giants’ Coronae,” 1996, Proposal selected for SAX Core Program.

[10] F. Reale, S. Sciortino, G. H. J. van den Oord & S. Serio, “The Coronae of Eclipsing RS CVn stars,” 1996, Proposal selected for SAX Core Program.

[11] G. Peres, F. Damiani, L. Golub & S. Saar, “Coronae of very late-type active stars,” 1996, Proposal selected for SAX Core Program.

[12] F. Bocchino, S. Sciortino, A. Maggio & O. Vancura, “The nature of the shell-type SNR X-ray Emission,” 1996, Proposal selected for SAX Core Program.

[13] R. Pallavicini, L. Pasquini & P. Molaro, “Lithium in turnoff stars of the old metal poor open cluster NGC2243,” 1996, Proposal selected for the ESO 3.6m telescope.

[14] S. Randich, R. Pallavicini & J.H.M.M. Schmitt, “The evolution of lithium abundance: K dwarfs in NGC6475,” 1996, Proposal selected for the ESO 3.6m telescope .

[15] S. Randich, R. Pallavicini, L. Pasquini & G. Tagliaferri, “Lithium and Oxygen Abundances in two intermediate age clusters,” 1996, Proposal selected for the ESO 3.6m telescope.

[16] A. Brown, M. Gagne’, R. Pallavicini, S. Drake, T. Ayres, G. Umana, C. Trigilio & K. Jones, “Linking the Impulsive and gradual phases of RS CVn flares: X-ray (SAX, XTE), EUV (EUVE), radio (VLA, AT, Noto), UV (IUE), and optical variability of HR1099,” 1996, Proposal accepted for the GO time on SAX (priority B).

[17] A. Maggio, S. Sciortino & A. Collura, “X-Ray Activity of Late-Type Giants in Bynary Systems,” 1996, Proposal selected by BMFT for ROSAT AO7 Guest Observer Program.

[18] S. Sciortino, F. Damiani, F. Favata, M. Lattanzi & J. Pye, “A Follow-up Deep Survey of the Young Open Cluster Stock 2,” 1996, Proposal selected by NASA for ROSAT AO7 Guest Observer Program.

[19] F. Damiani, G. Micela & S. Sciortino, “X-Ray Study of the Pre-Main-Sequence Population in Molecular Cloud,” 1996, Proposal selected by NASA for ROSAT AO7 Guest Observer Program.

[20] G. Micela, F. Favata, R. Jeffreys, F. Morale & J. Pye, “High Spatial Resolution Survey of the Young Open Cluster NGC 2516,” 1996, Proposal selected by SERC for ROSAT AO7 Guest Observer Program.

[21] F. Damiani, G. Micela & S. Sciortino, “High-Resolution X-ray Observations of Herbig BE Stars,” 1996, Proposal selected by NASA for ROSAT AO7 Guest Observer Program.

[22] R. Pallavicini, G. Micela, F. Morale, J. Pye, S. Sciortino, R. Garcia-Lopez & E. Martin, “A Follow-up Deep Survey of the Young Open Cluster Zeta Sculptoris,” 1996, Proposal selected by BMFT for ROSAT AO7 Guest Observer Program.

[23] F. Favata, G. Micela, S. Sciortino, J. Pye & R.S. Warwik, “The Rosat Pointed Galactic Plane Survey: Identification and Study of the Low Galactic Latitude Soft X-ray Population,” 1996, Proposal selected fot the 4 m Kurtis Schmidt Telescope.

[24] F. Favata, G. Micela & S. Sciortino, “Determination of the Initial Mass Function for low mass stars in the Sco-Cen star formation region,” 1996, Proposal selected for the ESO 2.2m telescope.

[25] F. Favata, G. Micela & S. Sciortino, “Determination of Accurate Atmospheric Parameters and Abundances in a Sample of Active Binaries,” 1996, Proposal selected for the ESO CAT telescope.

[26] G. Peres, A. Ciaravella, S. Orlando, F. Reale, R. Betta & S. Serio, “UVCS observations of the polar high speed solar wind generation region and CDS, EIT, and SUMER ancillary observations of the related chromospheric and coronal regions.,” 1996, Proposal selected for SOHO Observing Program.

[27] F. Damiani, G. Micela & S. Sciortino, “X-ray Studies of the Pre-Main-Sequence Stellar Population in Molecular Clouds,” 1995, Proposal selected by BMFT for ROSAT AO6 Guest Observer Program.

[28] A. Maggio, S. Sciortino & A. Collura, “Stellar X-ray Activity Near The Base Of The Giant Branch: Observations OF G and K Giants in Binary Systems,” 1995, Proposal selected by BMFT for ROSAT AO6 Guest Observer Program.

[29] S. Sciortino, F. Damiani, F. Favata, M. Lattanzi & J. Pye, “A Deep Survey of the Young Open Cluster Stock-2,” 1995, Proposal selected by NASA for ROSAT AO6 Guest Observer Program.

[30] G. Micela, F. Morale, R. Pallavicini & J. Pye, “A Deep Survey of the Young Open Cluster Zeta Sculptoris,” 1995, Proposal selected by SERC for ROSAT AO6 Guest Observer Program.

[31] F. Favata, S. Sciortino, G. Micela & F.R. Harnden,Jr., “A Deep Survey of the Open Cluster NGC 7092,” 1995, Proposal selected by BMFT for ROSAT AO6 Guest Observer Program.

[32] F. Bocchino, S. Sciortino, A. Maggio, F. Favata & J.C. Raymond, “Testing the Emission Conditions of the Vela Supernova Remnant,” 1995, Proposal accepted by ESO for optical Spectrophotometry at the 2.2m, La Silla, Chile.

[33] F. Bocchino, S. Sciortino, A. Maggio & J.C. Raymond, “Testing the Emission Conditions of the Vela Supernova Remnant,” 1995, Proposal accepted by ESA for UV Spectrophotometry with IUE.

[34] S. Serio, A. Maggio, G. Peres, F. Damiani & M. Itoh, “The Thermal Structure of dMe Stars’ Coronae,” 1995, Proposal for ASCA Observation (AO-4).

[35] F. Favata, N. Brickhouse, A.K. Dupree & A. Maggio, “Simultaneous ASCA and EUVE Observations of 44 Boo,” 1995, Accepted Proposal for ASCA Observation (AO-4).

[36] F. Damiani, G. Micela, S. Sciortino, V. Kashyap & R. Rosner, “A study of optical counterparts of ROSAT X-ray sources in star forming regions,” 1995, Proposal selected for the ARC telescope.

[37] F. Favata, G. Micela, S. Sciortino, V. Kashyap & R. Rosner, “Sloan photometry and spectroscopy of young solar-type stars in the NGC 2422 cluster,” 1995, Proposal selected for the ARC telescope.

[38] F. Favata, N. Trams, M. Lattanzi, G. Micela & S. Sciortino, “Lithium abundance, activity and rotation in the Stock 2 open cluster,” 1995, Proposal selected for the WHT telescope.

[39] F. Favata, G. Micela & S. Sciortino, “The lithium abundance of metalpoor K dwarfs.,” 1995, Proposal selected for the ESO CAT telescope.

[40] F. Favata, G. Micela, S. Sciortino & C. Corbally, “Photometric period determination of solar-type stars in the Stock 2 open cluster.,” 1995, Proposal selected for the 1.8m VATT telescope.

[41] R. Pallavicini, A. Ortolani, G. Tagliaferri & Y. Uchida, “Temperature and Abundance Diagnostics of the RS CVn Binaries Lambda Andromedae and Ar Piscium,” 1995, Proposal selected for ASCA (AO-4) observation program.

[42] M. Barbera, F. Favata, F. R. Harnden, Jr., G. Micela & S. Sciortino, “The Young Open Cluster NGC2422: Relation among Lithium Abundance, Rotation and Coronal Emission,” 1994, Proposal selected by NASA for ROSAT Guest Observer Program.

[43] F. Favata, M. Barbera, G. Micela & S. Sciortino, “A Deep HRI Observation of a WTTS-Rich Field in the Sco-Cen Star-Forming Region.,” 1994, Proposal selected by BMFT for ROSAT Guest Observer Program.

[44] J. F. Sckachter, F. R. Harnden, Jr., M. Barbera, G. Micela, S. Sciortino & F. Favata, “Identification of Selected Southern Einstein IPC Slew Survey Sources.,” 1994, Proposal selected by NASA for ROSAT Guest Observer Program.

[45] F. Damiani, S. Sciortino & G. Micela, “An X-ray Study of Stellar Formation Region Associated with Herbig Ae/Be Stars,” 1994, Proposal selected by BMFT for ROSAT Guest Observer Program.

[46] S. Sciortino, F. Damiani & G. Micela, “A Study of the X-ray Activity Rotation Relation for T-Tauri Stars,” 1994, Proposal selected by BMFT for ROSAT Guest Observer Program.

[47] F. R. Harnden, Jr., J. P. Caillault, G. Micela, C. F. Prosser, R. Rosner, S. Sciortino & J. R. Stauffer, “Completing the HRI observations of the Pleiades,” 1994, Proposal selected by NASA for ROSAT Guest Observer Program.

[48] G. Micela, J. Pye, F. Favata & S. Sciortino, “A Follow-up Variability Study of the serendipitously Discovered Very Active Nearby dM3 star G102-21.,” 1994, Proposal selected by SERC for ROSAT Guest Observer Program.

[49] F. Favata, G. Micela & S. Sciortino, “The Boron Abundance of Tidally Locked Binaries in the Hyades Cluster ,” 1994, Proposal selected by NASA for HST Guest Observer Program.

[50] S. Sciortino, F. Damiani & G. Micela, “The Origin of X-ray Emission of Herbig Ae/Be Stars,” 1994, Proposal selected for ASTROD Guest Observer Program.

[51] F. Damiani, S. Sciortino & G. Micela, “Observation of Rotational Modulation of X-ray Emission: Probing the Spatial Structure of T Tauri Stars Coronae,” 1993, Proposal selected by BMFT for ROSAT Guest Observer Program.

[52] F. R. Harnden, Jr., J. P. Caillault, G. Micela, C. F. Prosser, R. Rosner, S. Sciortino & J. Stauffer, “HRI Observations of the Pleiades,” 1993, Proposal selected by NASA for ROSAT Guest Observer Program.

[53] F. R. Harnden, Jr., J. Schachter, M. Barbera, G. Micela, S. Sciortino & F. Favata, “Identification of selected, southern Einstein IPC Slew-survey Sources,” 1993, Proposal selected by NASA for ROSAT Guest Observer Program.

[54] G. Micela, J. Pye, S. Sciortino & M. Barbera, “X-ray Coronal Spectra of Nearby Low-mass Old-Disk and Halo Populations Stars. A Follow Up Study.,” 1993, Proposal selected by SERC for ROSAT Guest Observer Program.

[55] S. Sciortino, A. Maggio, A. Collura & G. Micela, “Coronal Activity near the Base of the Red Giant Branch: Observations of G and K Giants in Binary Systems,” 1993, Proposal selected by BMFT for ROSAT Guest Observer Program.

[56] S. Sciortino, F. Damiani, G. Micela & J. Bouvier, “A Study of the X-ray Activity-Rotation Relation for T Tauri Stars,” 1993, Proposal selected by BMFT for ROSAT Guest Observer Program.

[57] G. Chincarini, O. Citterio, P. Conconi, T. Maccacaro, L. Stella, G. Tagliaferri, G. Trinchieri, A. Wolter, G. F. Bignami, F. Bocchino, A. Maggio, G. Micela, S. Sciortino, S. Serio, A. Collura, P. Giommi, L. Maraschi, R. Pallavicini, S. Pellegrini & G. Peres, “A 15-30 arcsec Resolution Replica X-ray Optics fo AXAF-S,” 1993, Proposal submitted to NASA and ASI.

[58] A. Maggio, F. Damiani, G. Peres, S. Sciortino & G. Micela, “Age Dependence of EUV Emission in Solar-Type Stars,” 1993, Proposal selected by NASA for EUVE Guest Observer Program.

[59] S. Sciortino & F. Bocchino, “Some Considerations on Scientific Impact of Proposed Italian Replica Mirrors for AXAF-S Mission: The case of SNR’s. ,” 1993, Presented at the March, 30, 1993 meeting of the AXAF-S team, in Milan.

[60] G. Micela, A. Maggio, R. Pallavicini, G. Peres, S. Sciortino & G. Tagliaferri, “Contributions to a proposal for AXAF-S: Normal Stars,” 1993, Presented at the March, 30, 1993 meeting of the AXAF-S team, in Milan.

[61] F. Favata, M. Barbera, G. Micela & S. Sciortino, “Lithium Abundance Determination in a Sample of Volume Limited Main-Sequence K Stars,” 1993, Observation Program Accepted at ESO for November 1993.

[62] F. Favata, M. Barbera, G. Micela, S. Sciortino & J. Schachter, “Identification and Classification of Candidate Stellar Einstein Slew Survey Sources,” 1993, Observation Program Accepted at ESO for January 1994.

[63] F. Favata, G. Micela & S. Sciortino, “Lithium Abundance Determination in a Sample of Active Binaries with Known Levels of X-ray Emission,” 1993, Observation Program Accepted at ESO for January 1994.